Paul's dad is five times as old as Paul. In 3 years, Paul will be 1/4 his dad's age. What are their ages now?
Answer by stanbon(75887)

Paul's dad is five times as old as Paul. In 3 years, Paul will be 1/4 his dad's age. What are their ages now?
-----------------------
Ages NOW DATA:
Paul's age = x ; Paul dad's age = 5x
--------------------------------------
Ages 3 years from NOW:
Paul's age = x+3 ; Dad's age = 5x+3
------------------------
EQUATION:
Paul age = (1/4)Dad's age using 3 yrs. from now data.
x+3 + (1/4)(5x+3)
4x+12 = 5x+3
x=9
Paul's age now is 9
Dad's age now is 45
Cheers,
Stan H.